# EMLoC: Emulator-based Memory-efficient Fine-tuning with LoRA Correction

### v1

@@ 5. $W^{c}_{B}\leftarrow W^{\prime}_{B}-\text{clamp}(\Delta,\lambda)$
Offsite [40] 84.3 91.3 69.0 77.7 51.0 76.1 45.9
UPop [32] 84.4 92.0 69.7 78.5 50.7 76.4 42.1
@@
Offsite [40] 84.0 92.3 69.6 77.2 50.6 76.6 45.4
UPop [32] 84.3 92.2 69.6 78.2 50.9 76.6 44.1
@@
Offsite [40] 83.9 91.8 67.0 77.1 50.1 75.2 44.2
UPop [32] 84.2 91.9 67.1 78.1 50.3 75.8 43.9
@@ 4.1 Experiment Setup
In our main experiments, we evaluate EMLoC on seven visual question answering (VQA) datasets. Four of these are widely used benchmarks: ChartQA [22], DocVQA [23], InfoVQA [24], and TextVQA [33]. To reflect realistic deployment scenarios that require domain adaptation, we also include three datasets from specialized domains: PMC-VQA [47] (medical imaging), WebSRC [5] (web page understanding), and WC-VQA [39] (multicultural knowledge). Ablation studies and additional analyses are primarily conducted on these three domain-specific datasets. For language modeling, we evaluate on two challenging mathematical reasoning benchmarks: MathQA [1] and GSM8K [7]. All reported metrics across datasets are based on accuracy.
@@
For the main experiments, we compare EMLoC against four baselines. Original: We directly fine-tune the original large model, InternVL2.5-8B [6], with LoRA. This baseline incurs significantly higher memory cost compared to other methods and serves as an upper bound in performance. Small: We fine-tune smaller variants of InternVL (4B and 2B), representing the low-resource alternative illustrated in Fig. 1(a). Offsite: We follow the Offsite-tuning [40] framework using the variant that omits full-model distillation to ensure fair comparison under the same compute constraints. UPop: A row-pruning-based method [32] is used to construct the emulator, followed by LoRA transfer using the procedure described in LORAM [45]. Except for the Small baseline, all methods—including ours—perform inference using the original InternVL2.5-8B model.
@@
Table 3: Applying EMLoC to NLP tasks. Following LORAM [45], a LLaMA2 13B is used as the full model. All experiments are conducted under the same budget.
@@
Table 4: Comparisons of emulator construction settings. Note that, LORAM [45] requires external data and additional continual pretraining for memory efficient finetuning. We validate that both row pruning and SVD can be applied for our emulator construction with supperior performances on various tasks using only downstream data.
@@
4.2 Main Results
Main Results.
We evaluate EMLoC using InternVL2.5-8B as the base model under three realistic memory-constrained fine-tuning scenarios. These settings simulate user environments with training memory budgets equivalent to: fine-tuning a 4B model (denoted as $50\%$ compression ratio in the table), fine-tuning a 2B model ($25\%$), and fine-tuning a 2B model using QLoRA [9] quantization ($25\%$+QLoRA). As shown in Table 1, EMLoC consistently outperforms all baselines across both standard and domain-specific VQA benchmarks. The improvements are especially significant on specialized datasets, where adaptation to the target domain is essential. Notably, EMLoC achieves performance closest to the upper bound set by full-model fine-tuning.
These results also underscore the adaptability of our approach under varying memory constraints, maintaining strong performance even under the tightest budgets. In contrast to the common strategy of fine-tuning smaller models, EMLoC provides a more effective alternative by better leveraging the capacity of the full model. In addition, comparisons with UPop and Offsite-tuning demonstrate the advantages of our downstream-aware emulator construction. Unlike row-pruning or layer-dropping approaches that restrict architectural flexibility, our method tailors the emulator to downstream data without altering model structure, enabling more effective and generalizable adaptation.
@@ 4.3 Analysis
We further evaluate EMLoC on larger-scale models, including the 26B and 38B variants of InternVL2.5. For this experiment, we construct a 5B-sized emulator, making the fine-tuning cost roughly equivalent to training a 5B model. We use LoRA rank 6 for the 26B model and rank 4 for the 38B model. As shown in Table 3, EMLoC consistently improves upon the zero-shot performance of these large models, demonstrating its effectiveness even at greater model scales. Importantly, all fine-tuning were conducted with less than 24GB of GPU memory—substantially lower than the 95GB typically required for directly fine-tuning the 38B model. This means that users capable of running inference can also perform fine-tuning using our method, without additional hardware requirements. These results underscore the practicality and scalability of EMLoC for resource-constrained users.
@@
We compare two emulator construction settings. The first is our proposed downstream-aware construction, in which the emulator is derived directly from task-specific data without any additional training. The second follows the approach adopted in LORAM, where the emulator is first constructed using general-purpose data, followed by continual pretraining. For general data, we use Laion-COCO [17]—a commonly used dataset for vision-language pretraining—and perform continual pretraining using 8 V100 GPUs.
As shown in Table 4, when row pruning is used as the emulator construction method, both settings achieve comparable downstream performance. However, the continual pretraining stage required in LORAM incurs substantial computational overhead, making it impractical for individual users. In contrast, when SVD is used for emulator construction, our downstream-aware method not only eliminates the need for continual pretraining but also achieves superior performance. Moreover, it incurs the lowest construction overhead, demonstrating the efficiency and effectiveness of our downstream-aware SVD.
Table 5: Ablation study on activation-aware SVD and LoRA correction.
Activation-Aware SVD LoRA Correction PMC-VQA WebSRC WC-VQA
51.0 74.4 44.7
51.5 79.0 45.8
51.2 74.4 44.8
51.6 79.6 46.2
@@
To evaluate the generality of our approach beyond vision-language tasks, we apply EMLoC to two challenging NLP datasets, MATHQA and GSM8K. We follow the experimental setup of LORAM and use LLaMA 2 13B [36] as the base model. As shown in Table 3, our method achieves comparable performance on MATHQA and outperforms LORAM on GSM8K, demonstrating its effectiveness in language-only settings. These results confirm that EMLoC is not limited to multimodal models and can be readily extended to other modalities. Furthermore, our method retains its core advantages—minimal memory overhead and no need for continual pretraining—making it more accessible to users with limited computational resources.
@@
Table 6: Performance comparisons of EMLoC before and after transferring the LoRA modules from the emulator to the original model. Note that running inference using emulator does not fully utilize the capacity of the original model and thus is not desirable.
@@
To better understand the role of the emulator and the effectiveness of our LoRA correction strategy, we compare model performance before and after transferring the fine-tuned LoRA modules from the emulator to the original model. As shown in Table 6, performance on the emulator itself is significantly lower the final model after LoRA transfer. This gap underscores the importance of using the full model for inference and justifies our design choice of transferring LoRA back.
@@
Figure 4: Sensitivity analysis of $\lambda$ in the LoRA correction algorithm. We plot performance on WC-VQA under different $\lambda$ values.
@@
We conduct ablation experiments to evaluate the contribution of two key components in the EMLoC framework: activation-aware SVD for downstream-aware emulator construction, and the LoRA correction algorithm for mitigating training-inference misalignment. As shown in Table 5, removing either component results in a drop in performance. This highlights the importance of preserving downstream-relevant knowledge during emulator construction, and the necessity of correcting for discrepancies introduced by training and inference on different models.
@@
We investigate the effect of the hyperparameter $\lambda$, which constrains the norm of the correction term $\Delta$ in our LoRA correction algorithm. As shown in Fig. 4, performance varies with different choices of $\lambda$. In particular, when no constraint is applied (i.e., $\lambda\rightarrow\infty$), performance degrades noticeably. We observe that, in such cases, the correction term may have a significantly larger norm than the original LoRA output. We hypothesize that correction term with large norm will dominate and distort the LoRA modules, ultimately harming downstream performance. In contrast, applying an appropriate constraint on the correction magnitude helps preserve the learned knowledge in the LoRA modules while still compensating for model misalignment.

### v2

@@ 4 Experiment
Offsite [xiao2023offsite] 84.3 91.3 69.0 77.7 51.0 76.1 45.9
UPop [pmlr-v202-shi23e] 84.4 92.0 69.7 78.5 50.7 76.4 42.1
@@
Offsite [xiao2023offsite] 84.0 92.3 69.6 77.2 50.6 76.6 45.4
UPop [pmlr-v202-shi23e] 84.3 92.2 69.6 78.2 50.9 76.6 44.1
@@
Offsite [xiao2023offsite] 83.9 91.8 67.0 77.1 50.1 75.2 44.2
UPop [pmlr-v202-shi23e] 84.2 91.9 67.1 78.1 50.3 75.8 43.9
@@ 4.1 Experiment Setup
In our main experiments, we evaluate EMLoC on seven visual question answering (VQA) datasets. Four of these are widely used benchmarks: ChartQA [masry-etal-2022-chartqa], DocVQA [mathew2021docvqa], InfoVQA [mathew2022infographicvqa], and TextVQA [singh2019towards]. To reflect realistic deployment scenarios that require domain adaptation, we also include three datasets from specialized domains: PMC-VQA [zhang2023pmc] (medical imaging), WebSRC [chen-etal-2021-websrc] (web page understanding), and WC-VQA [winata2024worldcuisines] (multicultural knowledge). Ablation studies and additional analyses are primarily conducted on these three domain-specific datasets. For language modeling, we evaluate on two challenging mathematical reasoning benchmarks: MathQA [amini-etal-2019-mathqa] and GSM8K [cobbe2021training]. All reported metrics across datasets are based on accuracy.
@@
For the main experiments, we compare EMLoC against four baselines. Original: We directly fine-tune the original large model, InternVL2.5-8B [chen2024expanding], with LoRA. This baseline incurs significantly higher memory cost compared to other methods and serves as an upper bound in performance. Small: We fine-tune smaller variants of InternVL (4B and 2B) with LoRA, representing the low-resource alternative illustrated in Fig.˜1(a). Offsite: We follow the Offsite-tuning [xiao2023offsite] framework using the variant that omits full-model distillation to ensure fair comparison under the same compute constraints. UPop: A row-pruning-based method [pmlr-v202-shi23e] is used to construct the emulator, followed by LoRA transfer using the procedure described in LORAM [zhang2025train]. Except for the Small baseline, all methods—including ours—perform inference using the original InternVL2.5-8B model.
@@
4.2 Main Results
Main Results.
We evaluate EMLoC using InternVL2.5-8B as the base model under three realistic memory-constrained fine-tuning scenarios. These settings simulate user environments with training memory budgets equivalent to: fine-tuning a 4B model (denoted as $50\%$ compression ratio in the table), fine-tuning a 2B model ($25\%$), and fine-tuning a 2B model using QLoRA [dettmers2023qlora] quantization ($25\%$+QLoRA). As shown in Table˜1, EMLoC consistently outperforms all baselines across both standard and domain-specific VQA benchmarks. The improvements are especially significant on specialized datasets, where adaptation to the target domain is essential. Notably, EMLoC achieves performance closest to the upper bound set by full-model fine-tuning.
These results underscore the adaptability of our approach under varying memory constraints, maintaining strong performance even under the tightest budgets. In contrast to the common strategy of fine-tuning smaller models, EMLoC provides a more effective alternative by better leveraging the capacity of the full model. In addition, comparisons with UPop and Offsite-tuning demonstrate the advantages of our downstream-aware emulator construction. Unlike row-pruning or layer-dropping approaches that restrict architectural flexibility, our method tailors the emulator to downstream data without altering model structure, enabling more effective and generalizable adaptation. Finally, we note that EMLoC is orthogonal with existing techniques, such as gradient checkpointing [chen2016training] and quantization [dettmers2023qlora]. In our experiments, EMLoC is used together with these techniques, yet it consistently outperforms other baselines. This demonstrates that EMLoC is compatible with existing memory reduction methods, and can be integrated with them to achieve further efficiency gains.
@@ 4.2 Main Results
Table 3: Applying EMLoC to NLP tasks. All settings follow LORAM [zhang2025train], and a LLaMA2 13B is used as the full model.
@@
Table 4: Comparisons of emulator construction settings. The first row corresponding to LORAM [zhang2025train] setting and the last row corresponding to EMLoC setting. Note that, LORAM requires external data and additional continual pretraining for memory efficient finetuning. We validate that both row pruning and SVD can be applied for our setting with better performances on various tasks using only a small subset of downstream data.
@@ 4.3 Analysis
We further evaluate EMLoC on larger-scale models, including the 26B and 38B variants of InternVL2.5. For this experiment, we construct a 5B-sized emulator, making the fine-tuning cost roughly equivalent to training a 5B model. We use LoRA rank 6 for the 26B model and rank 4 for the 38B model. As shown in Table˜3, EMLoC consistently improves upon the zero-shot performance of these large models, demonstrating its effectiveness even at greater model scales. Importantly, all fine-tuning were conducted with less than 24GB of GPU memory—substantially lower than the 95GB typically required for directly fine-tuning the 38B model. This means that users capable of running inference can also perform fine-tuning using our method, without additional hardware requirements. These results underscore the practicality and scalability of EMLoC for resource-constrained users.
@@
We compare two emulator construction settings. The first is our proposed downstream-aware construction, in which the emulator is derived directly from task-specific data without any additional training. The second follows the approach adopted in LORAM, where the emulator is first constructed using general-purpose data, followed by continual pretraining. For general data, we use Laion-COCO [laioncoco]—a commonly used dataset for vision-language pretraining—and perform continual pretraining using 8 V100 GPUs.
As shown in Table˜4, when row pruning is used as the emulator construction method, both settings achieve comparable downstream performance. However, the continual pretraining stage required in LORAM incurs substantial computational overhead, making it impractical for individual users. In contrast, when SVD is used for emulator construction, our downstream-aware method not only eliminates the need for continual pretraining but also achieves superior performance while incurring the lowest construction overhead, demonstrating the efficiency and effectiveness of our downstream-aware SVD. Furthermore, the performance gap observed in the last two rows highlights the importance of using downstream data rather than a general purpose data as the calibration set $\mathcal{D}_{C}$. By employing activation-aware SVD based on downstream activations, the constructed emulator better preserves task-relevant knowledge that facilitates subsequent fine-tuning.
@@
To evaluate the generality of our approach beyond vision-language tasks, we apply EMLoC to two challenging NLP datasets, MATHQA and GSM8K. We follow the experimental setup of LORAM and use LLaMA 2 13B [touvron2023llama] as the base model. As shown in Table˜3, our method achieves comparable performance on MATHQA and outperforms LORAM on GSM8K, demonstrating its effectiveness in language-only settings. These results confirm that EMLoC is not limited to multimodal models and can be readily extended to other modalities. Furthermore, our method retains its core advantages—minimal memory overhead and no need for continual pretraining—making it more accessible to users with limited computational resources.
@@
Table 5: Performance comparisons using finetuned emulators and EMLoC. Note with while both finetuned emulator and EMLoC share the same memory budget, the former does not fully utilize such budget during inference and thus is not desirable.
@@
To better understand the role of the emulator and the effectiveness of our LoRA correction strategy, we compare model performance before and after transferring the fine-tuned LoRA modules from the emulator to the original model. As shown in Table˜5, performance on the emulator itself is significantly lower the final model after LoRA transfer. This gap underscores the importance of using the full model for inference and justifies our design choice of transferring LoRA back.
@@
Table 6: Ablation study on activation-aware SVD and LoRA correction.
Ratio Activation-Aware SVD LoRA Correction PMC-VQA WebSRC WC-VQA
25% 51.0 74.4 44.7
51.2 74.4 44.8
51.5 79.0 45.8
51.6 79.6 46.2
50% 51.8 83.7 47.9
52.3 84.9 48.8
@@
We conduct ablation experiments to evaluate the contribution of two key components in the EMLoC framework: activation-aware SVD for downstream-aware emulator construction, and the LoRA correction algorithm for mitigating training-inference misalignment. As shown in Table˜6, removing either component results in a drop in performance. This highlights the importance of preserving downstream-relevant knowledge during emulator construction, and the necessity of correcting for discrepancies introduced by training and inference on different models.
By further examining these results, we observe that the effectiveness of LoRA correction seems to be closely related to the degree of misalignment between the emulator and the original model. When the emulator is constructed without activation-aware SVD, it fails to retain key downstream behaviors of the original model, leading to larger misalignment during fine-tuning and consequently weaker correction effects. Conversely, under milder compression settings where misalignment is reduced, the impact of LoRA correction becomes more pronounced. These findings suggest that while our current use of SVD-LLM [wang2025svdllm] provides a strong foundation for emulator construction, there remains room to explore more adaptive or task-aware compression strategies that could further enhance model alignment and correction effectiveness.
Applying to Diffusion Model Personalization
To further assess the generalizability of EMLoC beyond text generation, we apply it to the task of personalizing a large text-to-image diffusion model. Specifically, we use DreamBooth [ruiz2023dreambooth]—a well-established personalization method—combined with LoRA to fine-tune FLUX.1-dev [flux2024], a 12B state-of-the-art diffusion model. We conduct experiments on 8 subjects from the DreamBooth dataset, fine-tuning each for 500 steps without prior preservation. For evaluation, we report standard metrics used in prior works: DINO [caron2021emerging], CLIP-I [radford2021learning], and CLIP-T scores. Due to the short fine-tuning duration in this experiment, we adopt standard SVD for emulator construction instead of activation-aware SVD, which would be less cost-effective in this setting.
Quantitative results are presented in Table˜7. Note that normal fine-tuning requires over 35GB of memory, whereas EMLoC enables fine-tuning within a 24GB GPU budget. While EMLoC achieves slightly lower DINO and CLIP-I scores, it yields a higher CLIP-T score—likely due to slower convergence. We also show qualitative examples in Fig.˜6. These results demonstrate that EMLoC can be effectively applied to image generation tasks, reinforcing its potential as a plug-and-play replacement for conventional LoRA fine-tuning with reduced memory requirements.
Figure 4: Sensitivity analysis of $\lambda$ in the LoRA correction algorithm. We plot performance on WC-VQA under different $\lambda$.
Table 7: Quantitative results of applying EMLoC to diffusion model. EMLoC achieves comparable, sometimes slightly lower, performance than direct fine-tuning while significantly reducing memory usage.
Method
Fine-tuning
memory (GB)
DINO CLIP-I CLIP-T
w/o EMLoC 35.1 0.652 0.851 0.306
w/ EMLoC 22.9 0.615 0.831 0.321
@@
We investigate the effect of the hyperparameter $\lambda$, which constrains the norm of the correction term $\Delta$ in our LoRA correction algorithm. As shown in Fig.˜4, performance varies with different choices of $\lambda$. In particular, when no constraint is applied (i.e., $\lambda\rightarrow\infty$), performance degrades noticeably. We observe that, in such cases, the correction term may have a significantly larger norm than the original LoRA output. We hypothesize that correction term with large norm will dominate and distort the LoRA modules, ultimately harming downstream performance. In contrast, applying an appropriate constraint on the correction magnitude helps preserve the learned knowledge in the LoRA modules while still compensating for model misalignment.